Using the mean and the standard deviation you found in the second question​ (rounded to four​ decimals), answer the following probability questions using​ Stat>Calculators>Normal.

For​ each, copy and paste the StatCrunch output in the SHOW YOUR WORK.

Here is a short video that shows how to insert multiple graphs into the SHOW WORK​ area:

SHOW WORK

1. What is the probability that a randomly selected​ 5-year-old female will be taller than 40 inches​ tall? (Round to 4 decimal​ places)  

2. What is the probability that a randomly selected​ 5-year-old female will be shorter than 35 inches​ tall? (Round to 4 decimal​ places)

3. What is the probability that a randomly selected​ 5-year-old female will be between 39 and 47 inches​ tall? ​ (Round to 4 decimal​ places)

This is Question 2

44.5

42.4

42.2

46.2

45.7

44.8

43.3

39.5

45.4

43.0

43.4

44.7

38.6

41.6

50.2

46.9

39.6

44.7

36.5

42.7

40.6

47.5

48.4

37.5

45.5

43.3

41.2

40.5

44.4

42.6

42.0

40.3

42.0

42.2

38.5

43.6

40.6

45.0

40.7

36.3

44.5

37.6

42.2

40.3

48.5

41.6

41.7

38.9

39.5

43.6

41.3

38.8

41.9

40.3

42.1

41.9

42.3

44.6

40.5

37.4

44.5

40.7

38.2

42.6

44.0

35.9

43.7

48.1

38.7

46.0

43.4

44.6

37.7

34.6

42.4

42.7

47.0

42.8

39.9

42.3

Mean 42.2238

standard​ deviation 3.131
Answer #1

Given: u = 42.2238, = 3.131

We have, z = (x-u)/

1. What is the probability that a randomly selected​ 5-year-old female will be taller than 40 inches​ tall? (Round to 4 decimal​ places)  

= P(x > 40)

= P(z > (40-42.2238)/3.131)

= P(z > -0.710)

= 0.7611

2. What is the probability that a randomly selected​ 5-year-old female will be shorter than 35 inches​ tall? (Round to 4 decimal​ places)

= P(x < 35)

= P(z < (35-42.2238)/3.131)

= P(z < -2.307)

= 0.0105

3. What is the probability that a randomly selected​ 5-year-old female will be between 39 and 47 inches​ tall? ​ (Round to 4 decimal​ places)

= P(39 < x < 47)

= P((39-42.2238)/3.131 < z < (47-42.2238)/3.131)

= P(-1.030 < z < 1.525)

= P(z < 1.525) - P(z < -1.030)

= 0.9364 - 0.1515

= 0.7849